# AdvQuNN: A Methodology for Analyzing the Adversarial Robustness of Quanvolutional Neural Networks

### v1

@@ 4.1. Experimental Setup
In our experiment, we evaluated three machine learning models’ robustness against Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015), Projected Gradient Descent (PGD) (Madry et al., 2018), and Momentum-based Iterative Method (MIM) (Dong et al., 2018) attacks on MNIST (Deng, 2012) and Fashion MNIST (FMNIST) (Xiao et al., 2017) datasets: two classical models and one quantum model. The first classical model, depicted in Figure 4a, employs a convolutional layer with a $2\times 2$ kernel, stride of $2$, $4$ output channels, and ReLU activation, transforming $28\times 28$ pixel images into $14\times 14$ feature maps. These maps are then processed by a 10-neuron dense layer with softmax activation. In contrast, the second classical model relies solely on fully connected layers. The third model is a QuNN model with a similar structure to the first classical model for a fair comparison, as shown in Figure 4b, but incorporates a non-trainable quanvolutional layer with equivalent kernel size, stride, and output channels, yielding feature maps of analogous size from qubit expectation values, see Figure 2. The fully connected layers configuration in the DNN for MNIST follows a direct path to the yellow box as depicted in Figure 4. For FMNIST, additional layers are incorporated into the DNN for both classical and quantum models, as shown in Figure 4 red boxes.
@@
Figure 4. Diagram showing (a) a CNN with a convolutional layer, and (b) a QuNN with a quantum convolutional layer for FMNIST and MNIST datasets. Post-flattening, MNIST data follows a direct path (yellow box) to the softmax layer, while FMNIST data takes a red-box path with dense layers and dropout before softmax. Both paths converge for final model optimization and performance assessment.
@@
For each adversarial attack, we assess the robustness of the QuNNs for each Ansatz architecture in the quanvolutional filter (see 2). For each architecture, we run $7$ trials and plot the error bars chart. Each trial involves the following steps:
  1. (1)
    Quanvolve train and test images: Apply the quanvolutional layer with one type of Ansatz to the original $50$ train and $30$ test images, yielding quantum_train and quantum_test images.
  2. (2)
    Training DNN on quanvolved images: Training and validating the DNN model using quantum_train and quantum_test, with a batch size of $4$ and running the model for $30$ epochs.
  3. (3)
@@
    Create adversarial examples: Apply a type of attack on the $30$ original test images, followed by a quanvolutional layer, producing adversarial_quantum_test images. This is repeated for varied perturbation strength (epsilon) values. The values of epsilons are considering minor perturbations to relatively large perturbations to capture the behavior of the models on both spectrums.
  4. (4)
@@
    Evaluation: Finally we evaluate the trained models’ performance on the adversarial_quantum_test images across each epsilon value, resulting in accuracy versus epsilon plots.
@@
4.2. Comparative Analysis between CNNs and QuNNs against Adversarial Attacks
@@
We conduct experiments on three distinct models as discussed in Section 4.1, with a primary focus on the MNIST dataset. Here, we only focus on the ZZ full Ansatz (Figure 3b) to demonstrate key insights. Additionally, we conduct further analyses on the FMNIST dataset, highlighting its differences from the MNIST.
@@
In Figure 5a, subjected to FGSM attack, the classical models show a sharp drop in accuracy as $\epsilon$ increases, reaching near zero at smaller perturbations. In contrast, the quantum model with the ZZ full Ansatz retains higher accuracy at low perturbations (pointer 1⃝) and declines more gradually with increasing $\epsilon$ (pointer 2⃝), maintaining over $0.6$ accuracy even at higher values of $\epsilon$ (pointer 3⃝), indicating greater robustness compared to the classical models. The error bars in the quantum model indicate result variability, possibly due to quantum computing’s inherent randomness or other factors, which needs further investigation as discussed in Section 4.4.
Figure 5. Robustness evaluation of classical and quantum models with ZZ full Ansatz against a) FGSM, b) PGD, and c) MIM adversarial attacks at different perturbation strengths for the MNIST dataset.
@@
The figures show the three models’ robustness against FGSM, PGD, and MIM attacks, with the quantum model displaying superior accuracy under adversarial conditions on the MNIST dataset. For the FMNIST dataset, classical models’ accuracy quickly drops to zero at low epsilon values for all attacks. The quantum model also shows reduced accuracy, likely due to FMNIST’s greater complexity compared to MNIST. Despite a reduction in accuracy, the model still exhibits an overall trend of higher performance, maintaining an accuracy level around $0.4$, see Figures 6a, 6b, and 6c.
4.3. Comparative Analysis between Ansatz Architectures
Figure 6 illustrates the average accuracy of each architecture with error bars for each attack for the MNIST dataset. Each architecture shows different robustness against attacks as epsilon varies.
A) FGSM attack analysis, Figure 6d: As epsilon increases, the random architecture’s accuracy sharply drops, especially after the lowest perturbation level, yet shows moderate robustness at higher perturbations. The ZZ linear architecture initially outperforms the Random architecture but aligns with it at higher epsilons. The no-entanglement architecture declines more gradually, indicating better overall FGSM attack resilience. Both ZZ full and star architectures maintain high accuracy across all perturbation levels, showing the greatest robustness among the five architectures.
B) PGD attack analysis, Figure 6e: The random and ZZ linear architectures show a notable accuracy drop at low epsilon but stabilize at similar levels for high perturbations. The no-entanglement architecture, more stable than Random and ZZ linear, suggests better PGD attack resilience. Consistently with FGSM results, ZZ full and star architectures demonstrate the highest robustness under PGD attack as well.
C) MIM attack analysis, Figure 6f: The random architecture is consistently the least robust, showing the lowest accuracy across all perturbation levels. ZZ linear and no-entanglement architectures perform similarly, maintaining a slightly better hold on accuracy at high perturbations. ZZ full and star entanglement architectures again prove the most robust, with a slower decline and the highest accuracy across all perturbations.
For the MNIST dataset, ZZ full and star entanglement architectures are most robust against all three adversarial attacks, showing strong resilience. The no-entanglement and ZZ linear exhibit intermediate performance, with no-entanglement generally surpassing ZZ linear, especially in FGSM and PGD attacks. The Random architecture is the least robust, with a notable accuracy decline across all attacks, particularly at lower perturbations.
With the FMNIST dataset (Figures 6a, 6b, and 6c), the Random architecture outperforms others (pointer 2⃝, Figure6a), despite not achieving the highest accuracy at zero epsilon (pointer 1⃝, Figure 6a). Other architectures show consistent performance across varying epsilon levels and stabilize at high $\epsilon$ (e.g., pointer 1⃝, Figure 6b). Notably, the Ansatz architecture’s robustness ranking of the FMNIST differs from the MNIST, indicating that the Ansatz architecture design is greatly affected by the dataset it is applied to. This suggests a close link between the QuNN model’s robustness and its quantum circuit architecture, depending on the dataset.
@@ 4.3. Comparative Analysis between Ansatz Architectures
Figure 6. Evaluation of the robustness of the 5 Ansatz architectures under three adversarial attacks: a) FGSM, b) PGD, and c) MIM for the FMNIST dataset and d) FGSM, e) PGD, and f) MIM for the MNIST dataset.
4.4. Remarks and dicsussion
@@
The accuracy of all Ansatz architectures plateaus beyond a perturbation strength of $1$ (e.g. pointer 4⃝, Figure 5a), suggesting a QuNN robustness threshold at this level. Beyond this threshold, the accuracy does not further decline, maintaining a certain level rather than dropping to zero. This accuracy varies for each model training run and adversarial example generation, as indicated by the error bar lengths. Since FGSM, PGD, and MIM attacks use gradient information that varies across training sessions, the adversarialness of examples differs, affecting the interaction of the fixed quanvolutional layer with these images and resulting in variable error bar sizes.
@@
The non-trainable quanvolutional layer acts as a low-pass filter, which reduces extreme noise and retains some key features of the input image that the model was initially trained on, despite adversarial noise. This allows the model to still recognize patterns of the original class, preventing further accuracy drops. The plateau in performance likely indicates the model’s best attempt to classify severely perturbed images, demonstrating its robustness against adversarial attacks. If this quanvolutional filter is designed adequately, reflecting the architecture of the Ansatz, it might preserve certain structural features of the input images. For instance, for the MNIST dataset, a common characteristic between the $ZZ$ full and $ZZ$ star architectures, both entanglement patterns involve the first qubit getting entangled with all others. We can seek such characteristics to design tailored Ansatz architecture specific to the dataset.
5. Conclusion
This study focused on examining the security of QuNNs from a new perspective, not only by comparing them to classical models, but also by methodically assessing the impact of Ansatz design of the quanvolutional filter on the robustness of QuNNs to advanced attacks. In the context of a white-box attack environment, we conducted a thorough QuNN security evaluation considering five Ansatz architectures: No entanglement, ZZ full entanglement, ZZ linear entanglement, ZZ star entanglement, and Random architecture. Our analysis revealed that ZZ full and ZZ star entanglements significantly influence QuNNs’ robustness when applied to the MNIST dataset. Despite confirming recent claims about the robustness of QuNNs, this study goes a step further by demonstrating their relativity to the quantum circuit design, marking this study the first research to uncover these important insights.
This research demonstrates that QuNNs not only maintain high accuracy levels compared to traditional DNNs, but can also achieve enhanced accuracy with certain Ansatz architectures. We observed that QuNNs require more significant perturbations to experience a decrease in accuracy, whereas the classical models are susceptible to accuracy loss even with minor perturbations. Furthermore, our findings highlight that the architecture of the Ansatz in the quantum circuits of QuNNs is heavily influenced by the specific nature of the problem being addressed.
In our future research, we intend to extend our findings to a wider range of QML models, applying them across a broader spectrum of datasets and under diverse attack scenarios. We also plan to design specialized quantum circuits that are more data- and attack-oriented, with the goal of achieving enhanced robustness. Additionally, we will delve deeper into the accuracy trends of QuNNs and explore strategies to narrow the error margins, thereby minimizing variance in our results.
In conclusion, beyond the conventional methods for enhancing the robustness of QuNNs like adversarial training, this study paves new pathways for robustifying QuNN systems. By custom-designing quantum circuit architectures to specifically counter adversarial attacks, it is possible to embed inherent robustness directly into the model’s structure. This approach represents a significant shift in how quantum circuits can improve QML systems’ resilience against such threats.
Acknowledgements.

### v2

@@ IV-A Experimental Setup
In our experiment, we evaluated three machine learning models’ robustness against Fast Gradient Sign Method (FGSM) [28], Projected Gradient Descent (PGD) [29], and Momentum-based Iterative Method (MIM) [30] attacks on MNIST [31] and Fashion MNIST (FMNIST) [32] datasets: two classical models and one quantum model. The first classical model, depicted in Figure 4a, employs a convolutional layer with a $2\times 2$ kernel, stride of $2$, $4$ output channels, and ReLU activation, transforming $28\times 28$ pixel images into $14\times 14$ feature maps. These maps are then processed by a 10-neuron dense layer with softmax activation. In contrast, the second classical model relies solely on fully connected layers. The third model is the QuNN model with a similar structure to the first classical model for a fair comparison, Figure 4b. It includes a non-trainable quanvolutional layer with the same kernel size, stride, and output channels, yielding feature maps of analogous size from qubit expectation values, see Figure 2. The fully connected layers configuration in the DNN for MNIST follows a direct path to the yellow box as depicted in Figure 4. For FMNIST, additional layers are incorporated into the DNN for both classical and quantum models, as shown in Figure 4 red boxes.
@@
Figure 4: Diagram showing (a) a CNN with a convolutional layer, and (b) a QuNN with a quantum convolutional layer for FMNIST and MNIST datasets. Post-flattening, MNIST data follows a direct path (yellow box) to the softmax layer, while FMNIST data takes a red-box path with dense layers and dropout before softmax. Both paths converge for final model optimization and performance assessment.
@@
For each adversarial attack, we assess the robustness of the QuNNs for each Ansatz architecture in the quanvolutional filter (see Figure 2). For each architecture, we run $7$ trials and plot the error bars chart. Each trial involves the following steps:
  1. 1.
    Train the QuNN model: We train and validate the QuNN model using $50$ training samples. Each image is processed through a quanvolutional layer, followed by a classical layer for classification. The model is trained with a batch size of $4$ for $30$ epochs, using a learning rate of $0.001$. The classical layer parameters are updated iteratively.
  2. 2.
@@
    Create adversarial examples: Using the trained model we apply white box FGSM, PGD, and MIM attack on $30$ test images using Algorithms 1, 2, and 3 to create adversarial test images. The perturbations in the image pixels are mapped to rotational perturbations on qubits. This is repeated for varied perturbation strength (epsilon) values. The values of epsilons are considering minor perturbations to relatively large perturbations to capture the behavior of the models on both spectrums.
  3. 3.
@@
    Evaluation: Finally we evaluate the performance of the trained QuNN on adversarial test images at various epsilon values, resulting in accuracy versus epsilon plots.
@@
IV-B Comparative Analysis between CNNs and QuNNs against Adversarial Attacks
@@
We conduct experiments on three distinct models as discussed in Section IV-A, with a primary focus on the MNIST dataset. Here, we only focus on the ZZ full Ansatz (Figure 3b) to demonstrate key insights. Additionally, we conduct further analyses on the FMNIST dataset, highlighting its differences from the MNIST.
@@
In Figure 5a, subjected to FGSM attack, the classical models show a sharp drop in accuracy as $\epsilon$ increases, reaching near zero at smaller perturbations. In contrast, the quantum model with the ZZ full Ansatz retains higher accuracy at low perturbations (pointer ①) and declines more gradually with increasing $\epsilon$ (pointer ②), maintaining over $0.6$ accuracy even at higher values of $\epsilon$ (pointer ③), indicating greater robustness compared to the classical models. The error bars in the quantum model indicate result variability, possibly due to quantum computing’s inherent randomness or other factors as discussed in Section IV-D.
Figure 5: Robustness evaluation of classical and quantum models with ZZ full Ansatz against a) FGSM, b) PGD, and c) MIM adversarial attacks at different perturbation strengths for the MNIST dataset.
@@
For the FMNIST dataset, classical models’ accuracy quickly drops to zero at low epsilon values for all attacks. The quantum model also shows reduced accuracy, likely due to FMNIST’s greater complexity compared to MNIST. Despite a reduction in accuracy, the model still exhibits an overall trend of higher performance, maintaining an accuracy level around $0.4$, see Figures 6a, 6b, and 6c.
IV-C Comparative Analysis between Ansatz Architectures
@@
Figure 6: Evaluation of the robustness of the 5 Ansatz architectures under three adversarial attacks: a) FGSM, b) PGD, and c) MIM for the FMNIST dataset and d) FGSM, e) PGD, and f) MIM for the MNIST dataset.
In the analysis of adversarial attacks on various quantum circuit architectures, Figures 6d, 6e, and 6f, illustrate the average accuracy of each architecture with error bars for each attack, showing different robustness against adversarial attacks across the MNIST dataset as epsilon varies. The ZZ full and star entanglement architectures exhibit superior robustness compared to all architectures across all three types of attacks (FGSM, PGD, and MIM), maintaining high accuracy even at high perturbation levels. In contrast, the no-entanglement and ZZ linear architectures exhibit better resilience than the Random architecture, showing a more gradual decline in accuracy. On the other hand, the no-entanglement slightly surpasses the ZZ linear and behaves similarly for the MIM attack at high epsilons. Conversely, the Random architecture consistently underperforms, with a sharp accuracy decline at lower perturbations, stabilizing only at higher levels.
With the FMNIST dataset (Figures 6a, 6b, and 6c), the Random architecture outperforms others (pointer ②, Figure6a), despite not achieving the highest accuracy at zero epsilon (pointer ①, Figure 6a). Other architectures show consistent performance across varying epsilon levels and stabilize at high $\epsilon$ (e.g., pointer ①, Figure 6b). Interestingly, the robustness ranking of the Ansatz architecture varies between the FMNIST and MNIST datasets. This difference highlights the impact of the chosen dataset on the Ansatz architecture’s performance. It suggests a strong link between the robustness of the QuNN model and the configuration of its quantum circuit, which is dependent on the specific dataset being employed.
IV-D Remarks and discussion
The QuNN circuit includes angle encoding combined with a circuit that exhibits various entanglement scheme patterns. While angle encoding alone can increase robustness due to better decision boundary separation [19, 33], at low epsilons, this combination enhances the QuNN’s resilience against adversarial attacks as the circuit’s entanglement increases. This is evident in the results for the MNIST dataset, where the transition from no entanglement (Figure 3a) to full entanglement (Figure 3b) demonstrates improved resilience. Recent theoretical studies [25] have shown that angle encoding, which involves local operations, when combined with a quantum circuit exhibiting quantum chaos (high local-operator entanglement), can provide robustness against universal adversarial attacks. Our work provides empirical proof for this theoretical study.
@@ IV-D Remarks and discussion
All Ansatz architectures show a plateau in accuracy when perturbation strength exceeds $1$ (see pointer ④, Figure 5a). This plateau occurs because image pixel values are clipped between 0 and 1, which makes all attacked images look identical beyond epsilon equal to 1. Consequently, the QuNN performance is the same. This accuracy varies for each model training run and adversarial example generation, as indicated by the error bar lengths. Since FGSM, PGD, and MIM attacks use gradient information that varies across training sessions, the adversarialness of examples differs, affecting the interaction of the fixed quanvolutional layer with these images and resulting in variable error bar sizes.
@@
The non-trainable quanvolutional layer in QuNN acts as a low-pass filter, reducing extreme perturbations and retaining robust, meaningful features of the image, whereas classical models depend on informative but vulnerable features [22], which explains the significant difference in performance. As a result, QuNN’s performance only decreases significantly under very high perturbations. If this quanvolutional filter is designed adequately, reflecting the architecture of the Ansatz, it may enhance the extraction of robust features from the input images. For instance, for the MNIST dataset, a common characteristic between the $ZZ$ full and $ZZ$ star architectures, both entanglement patterns involve the first qubit getting entangled with all others. We can seek such characteristics to design tailored Ansatz architecture specific to the dataset.
V Conclusion
The AdvQuNN analysis reveals key properties that enhance QuNN model robustness against adversarial attacks. By examining various Ansatz architectures, it is evident that qubit connectivity patterns significantly influence data extraction and robustness. Specifically, for the MNIST dataset, increased qubit entanglement allows the model to detect and learn robust features, making it harder for adversaries to introduce effective perturbations. This confirms the experiment results where QuNNs require more significant perturbations to experience a decrease in accuracy, whereas the classical models are susceptible to accuracy loss even with minor perturbations. Despite confirming recent claims about the robustness of QuNNs, this study goes a step further by demonstrating their relativity to the quantum circuit design, marking this study the first research to uncover these important insights.
In future research, we aim to extend our findings to various QML models across more datasets and attack scenarios. We plan to design specialized, data- and attack-oriented quantum circuits to enhance robustness, delve into QuNN accuracy trends, and explore strategies to minimize error margins. This study paves new pathways for robustifying QuNN systems by custom-designing quantum circuit architectures to counter adversarial attacks, embedding inherent robustness directly into the model’s structure, representing a significant shift in improving QML resilience against threats.
Acknowledgment